if Qs = 500+50P and Qd = 125-25p find Q* and P*

1
Expert's answer
2021-09-26T20:14:20-0400

At equilibrium

"Q_s=Q_d \\

500+50P=125-25P \\

375= -75P \\

P^*=\frac{375}{-75} =-5 \\

Q^* = 500 + 50 \times (-5) = 250"